OpenStudy (anonymous):

You are performing a cohort study to look at risk factors for breast cancer in postmenopausal women. If the probability of developing breast cancer among smokers is .05 for the study duration, then if you sample 100 smokers, how many do you expect to develop the disease? Give the expected value +/- 1 standard deviation.

OpenStudy (tkhunny):

Have you considered the Binomial Approximation? Mean = n*p Variance = n*p*q